OpenStudy (lgbasallote):

Prove that the limit of sinx/x as x approaches 0 equals 1

OpenStudy (anonymous):

i can do the proof but.......

OpenStudy (anonymous):

look in any calc book and you will see it. annoying complicated for such a simple problem. uses "squeeze lemma" lol

OpenStudy (anonymous):

I guess L'Hopital rule would work

OpenStudy (anonymous):

0/0 so L'H Cos(x)/1 Cos(0)=1

OpenStudy (valpey):

Or use L'Hospital's rule: \[lim_{x\rightarrow 0}\frac{sin(x)}{x} = lim_{x\rightarrow 0}\frac{\frac{d(sin(x))}{dx}}{\frac{d(x)}{dx}} \] \[= lim_{x\rightarrow 0}\frac{cos(x)}{1} = \frac{1}{1} = 1\]